OpenStudy (anonymous):

8x+9=10x+2 what does it equal

OpenStudy (anonymous):

The left side equals the right side. That's why things with an EQUAL sign in them are called EQUATions. I think what you wanted to ask is "solve for x". And to figure that out, you want to get the numbers on one side, and the "x" values on the other. Do you know how to do that?

OpenStudy (anonymous):

bo i dont know how to solve it

OpenStudy (mertsj):

1. Subtract 8x from both sides 2. Subtract 2 from both sides. 3. Divide both sides by 2.

OpenStudy (anonymous):

\[8x+9=10x+2\] Gathering the terms on opposite sides, subtract 8x and 2 from both sides. \[(8x-8x)+(9-2)=(10x-8x)+(2-2)\] Simplifying, \[2x=7\] Dividing both sides by 2, get get \[x=3.5\]

OpenStudy (anonymous):

So your problem "equals" 37.